How many molecules of hydrogen gas are there in 8.32 g of Hepatanoic acid?


Molar Mass of the Acid = 130.182 g/mol

Molar Mass of Hydrogen = 1.008 g/mol


1
Expert's answer
2021-06-14T13:38:19-0400

Solution:

Heptanoic acid = C7H14O2

Hydrogen gas = H2


C7H14O2 → 7H2

According to the scheme above: n(C7H14O2) = n(H2)/7

Thus:

(8.32 g C7H14O2) × (1 mol C7H14O2 / 130.182 g C7H14O2) × (7 mol H2 / 1 mol C7H14O2) = 0.4474 mol H2


One mole of any substance contains 6.022×1023 atoms/molecules.

Therefore,

Number of H2 molecules = (0.4474 mol) × (6.022×1023 molecules / 1 mol) = 2.69×1023 H2 molecules


Answer: 2.69×1023 molecules of hydrogen gas (H2).